# Understanding Optimal Feature Transfer via a Fine-Grained Bias-Variance Analysis

### v1

@@ F.1 Existence and Uniqueness of $h_{0}$
Proof of Proposition F.1.
Recall from the statement of Proposition F.1 that $h_{1}>n$. Now let us state the following claim.
@@
for $\tilde{\mathcal{V}}$ defined in (59). Recall from Proposition E.2 and its proof that $\tilde{\mathcal{V}}=g\quantity(\sum_{i\in H}x_{i})$ where $g$ is a strictly increasing convex function on the range of $\sum_{i\in H}x_{i}$. It is therefore sufficient to consider the convex optimization problem of minimizing $\frac{1}{2h}\sum_{i\in H}x_{i}^{2}$ under the linear constraint $\frac{1}{h}\sum_{i\in H}x_{i}=1-\frac{n}{h}$. The corresponding Lagrangian is
@@
where $\rho\in\mathbb{R}$ is the Lagrange multiplier. Minimizing the Lagrangian yields the optimal solution $x_{i}=1-n/h,\forall i\in H$. The result then follows from Proposition E.1.
@@
for $h_{0}\in\{n,...,h_{1}\}$ defined in Proposition F.1, then the KKT conditions hold.
@@
where we used $\varphi_{i}\geq\varphi_{h_{0}},\forall i\in\{1,...,h_{0}\}$ in the first inequality and Proposition F.1 in the second inequality. This implies that $x_{i}\in[0,1],\forall i\in\{1,...,h_{0}\}$. We may then conclude that the primal feasibility condition holds.
@@
where we used $\varphi_{i}\leq\varphi_{h_{0}},\forall i\in\{h_{0}+1,...,h\}$ in the first inequality and Proposition F.1 in the second inequality. The above then implies that
@@
We have shown above that for $\quantity{x_{i}}_{i\in H}$ defined in (78) and KKT multipliers defined in (76), the KKT conditions hold. Note that the objective of (71) is differentiable and recall from Proposition E.2 that it is also convex. Furthermore, we have proved $\quantity{x_{i}}_{i\in H}$ defined in (78) satisfies the constraint of (71); this along with the fact that all constraints of (71) are linear implies that the Slater’s condition holds (see [10] Section 5.2.3 for a review of the Slater’s condition). Given the above, we know from [10] Section 5.5.3 that KKT conditions are sufficient for optimality. It follows that $\quantity{x_{i}}_{i\in H}$ defined in (78) is the optimal solution of (71). The optimal choice of $\quantity{\hat{r}_{i}}$ then follows Proposition E.1.
@@
Proposition F.2 below justifies the statement of Remark 5.2.
Proposition F.2 (Fully optimized $\widehat{\mathbf{B}}$ w.r.t. $\mathcal{V}$).
@@
Proof of Proposition F.2.
Recall from the proof of Lemma B.2 in Appendix B that given the change of variable
@@
This is precisely the optimization problem (71) we studied in the proof of Theorem 5.1 where we obtained the the optimal solution is $x_{i}=1-n/h,\forall i\in H$. However, since we no longer assumes (17), we cannot apply Proposition E.1. Instead, we observe the candidate solution
@@
Appendix G Obtain Fully-Optimized $\widehat{\mathbf{B}}$ via Backpropogation
@@
Backpropagating through Eigendeocmposition.
@@
Appendix H Upstream Sample Complexity.
In this section, we relax the assumption of oracle knowledge of the ground-truth representation $\mathbf{B}^{\star}$. There are many possible models for how these unknowns could be estimated from pretraining data, and our previous results are not tied to any particular setup. We consider the simple example given in Example 2.1 and track how these estimation errors affect the downstream risk. We recall the setting here for reader’s convenience. Recall that in the upstream, the model receives training data for $q$ distinct upstream tasks $\quantity(\mathbf{y}_{\mathsf{pre}}^{(i)},\mathbf{X}_{\mathsf{pre}}^{(i)})_{i=1}^{q}$ where
@@
Assume that the pretraining data are abundant $n_{\mathsf{pre}}>p$ and the OLS estimators $\tilde{\mathbf{B}}^{\star}=[\tilde{\mathbf{b}}^{(1)},...,\tilde{\mathbf{b}}^{(q)}],\tilde{\mathbf{b}}^{(i)}=\quantity({\mathbf{X}_{\mathsf{pre}}^{(i)}}^{\top}\mathbf{X}_{\mathsf{pre}}^{(i)})^{-1}{\mathbf{X}_{\mathsf{pre}}^{(i)}}^{\top}\mathbf{y}_{\mathsf{pre}}^{(i)}$ are used to estimate $\mathbf{B}^{\star}$. Let ~$\tilde{\mathfrak{R}}^{\mathsf{avg}}$ be the objective $\mathfrak{R}^{\mathsf{avg}}$ based on the inaccurate estimate $\tilde{\mathbf{B}}^{\star}$. The result below characterizes the error in approximating $R$ with ~$\tilde{\mathfrak{R}}^{\mathsf{avg}}$. See proof in Appendix H.
Theorem H.1.
@@
We prove Theorem H.1 below. Recall definition of $\tilde{\mathbf{B}}^{\star}$ from LABEL:SectionPrertain. Let us define the following notation
@@
We see from Lemma H.2 that $\mathfrak{B}-\tilde{\mathfrak{B}}$ depends on the quantity $\norm{(\tilde{\mathbf{B}}^{\star}-\mathbf{B}^{\star})\bm{\alpha}^{\star}}_{2}$. The following lemma bounds this quantity using Hanson-Wright inequality, which we state in Lemma C.1.
Lemma H.3.
@@
Now we are ready to prove the main result Theorem H.1.
Proof of Theorem H.1.
Combining Lemma H.2 and Lemma H.3 yields the following statement: conditioned on $\bm{\alpha}^{\star}$, for any $D>0$, there exists a constant $C(D)$ such that with probability at least $1-p^{-D}$,
@@
along with Lemma B.1 and Lemma B.2, we have the further upper bound that
@@
Now using the above, along with Lemma B.1 and Lemma B.2, we have the following statement: for any $D>0$, there exists a constant $C(D,M)$ such that with probability at least $1-p^{-D}$,
@@
Appendix I Supplementary Figures
I.1 Compare Fine-Grained Risk Components in Figure 2
@@
I.2 Figure 2 with different choices of $q$
@@
I.3 Ablating $q$, SNR and Common Structure in $\mathbf{B}^{\star}$
@@
I.4 Optimize $\mathcal{V},\mathfrak{B}$ alone
We reproduce Figure 3 under exact settings except that we optimizes for $\mathfrak{B}^{\mathsf{avg}}$ only. This gives us Figure 7, (a)-(f). We do the same but optimizes for $\mathcal{V}$ only. This gives us Figure 7, (g)-(l).
Observe that in the hard-selection regime (c), we have $50$ non-zero eigenvalues exactly equal to $q$ and they are all separate from 00 by a noticeable gap. Compare this with the soft-selection regime (f) where we have about $120$ non-zero eigenvalues less than $q=150$ and there is no separation from zero. Observe in (h), (k), the eigenvectors of $\widehat{\mathbf{B}}\widehat{\mathbf{B}}^{\top}$ align with eigenvectors of $\mathbf{\Sigma}$ in exactly the reverse order.

### v2

@@ 6.2.1 Existence and Uniqueness of $h_{0}$
Proof of Proposition 6.4.
Recall from the statement of Proposition 6.4 that $h_{1}>n$. Now let us state the following claim.
@@ 6.2.2 Analysis of KKT condition
for $\tilde{\mathcal{V}}$ defined in (74). Recall from Proposition D.2 and its proof that $\tilde{\mathcal{V}}=g\quantity(\sum_{i\in H}x_{i})$ where $g$ is a strictly increasing convex function on the range of $\sum_{i\in H}x_{i}$. It is therefore sufficient to consider the convex optimization problem of minimizing $\frac{1}{2h}\sum_{i\in H}x_{i}^{2}$ under the linear constraint $\frac{1}{h}\sum_{i\in H}x_{i}=1-\frac{n}{h}$. The corresponding Lagrangian is
@@
where $\rho\in\mathbb{R}$ is the Lagrange multiplier. Minimizing the Lagrangian yields the optimal solution $x_{i}=1-n/h,\forall i\in H$. The result then follows from Proposition D.1.
@@
for $h_{0}\in\{n,...,h_{1}\}$ defined in Proposition 6.4, then the KKT conditions hold.
@@
where we used $\varphi_{i}\geq\varphi_{h_{0}},\forall i\in\{1,...,h_{0}\}$ in the first inequality and Proposition 6.4 in the second inequality. This implies that $x_{i}\in[0,1],\forall i\in\{1,...,h_{0}\}$. We may then conclude that the primal feasibility condition holds.
@@
where we used $\varphi_{i}\leq\varphi_{h_{0}},\forall i\in\{h_{0}+1,...,h\}$ in the first inequality and Proposition 6.4 in the second inequality. The above then implies that
@@
We have shown above that for $\quantity{x_{i}}_{i\in H}$ defined in (60) and KKT multipliers defined in (58), the KKT conditions hold. Note that the objective of (53) is differentiable and recall from Proposition D.2 that it is also convex. Furthermore, we have proved $\quantity{x_{i}}_{i\in H}$ defined in (60) satisfies the constraint of (53); this along with the fact that all constraints of (53) are linear implies that the Slater’s condition holds (see [10] Section 5.2.3 for a review of the Slater’s condition). Given the above, we know from [10] Section 5.5.3 that KKT conditions are sufficient for optimality. It follows that $\quantity{x_{i}}_{i\in H}$ defined in (60) is the optimal solution of (53). The optimal choice of $\quantity{\hat{r}_{i}}$ then follows Proposition D.1.
@@
Proposition 6.5 below justifies the statement of Remark 5.2.
Proposition 6.5 (Fully optimized $\widehat{\mathbf{B}}$ w.r.t. $\mathcal{V}$).
@@
Proof of Proposition 6.5.
Recall from the proof of Lemma 6.2 in Section 6.1 that given the change of variable
@@
This is precisely the optimization problem (53) we studied in the proof of Theorem 5.1 where we obtained the the optimal solution is $x_{i}=1-n/h,\forall i\in H$. However, since we no longer assumes (14), we cannot apply Proposition D.1. Instead, we observe the candidate solution
@@
7 Discussion and Future Directions
Appendix A Proof of Proposition 2.2
Proof of Proposition 2.2.
where
It then follows from standard OLS formulas that
Note that $\hat{\bm{\Lambda}}$ contains matrix quantities
Appendix B Downstream Risk Concentration
Lemma B.1 (Hanson-Wright Inequality).
Let $\bm{\xi}$ be some random vector satisfying conditions in Lemma B.2 and $\mathbf{A}$ be any $n\times n$ real-valued matrix. There exists an absolute constant $c>0$ such that for any $x>0$, with probability larger than $1-\exp(-x)$, we have
where $\norm{\cdot}_{F}$ denotes Frobenius norm.
We first prove the following lemma using Hanson-Wright inequality.
Lemma B.2.
Proof of Lemma B.2.
Let us introduce the notation
Then, we can write
[…]
Using the elementary inequality $x/(1+xb_{0})\leq(4b_{0})^{-1}$, Lemma 6.1, Lemma 6.2 and 3.1, we obtain that
It then follows from 3.1 and the above that
we have that
The required statement follows if we set $x\leftarrow D\cdot\log q$. ∎
We are now ready to prove Proposition 4.3.
Proof of Proposition 4.3.
Using triangle inequality, we have that
[…]
Using the above, Lemma 6.1 and Lemma 6.2, we obtain
[…]
The result then follows from Theorem 3.3, Lemma B.2 and an application of the union bound.
Appendix C Minimax Optimality
where, with $\sigma_{\mathrm{max}}(\cdot)$ denoting the top eigenvalue,
Instead of tackling the bi-level optimization on the LHS, we may minimize RHS using backpropagation routine developed in Appendix E. For the latter, note that it is standard to differentiate through eigendecomposition and thus $\sigma_{\mathrm{max}}(\cdot)$. See Appendix E for more detailed discussion and pointers to relevant reference. Furthermore, as we will show in Appendix D, $\mathfrak{B}^{\mathsf{worst}}$ may be written into a convex objective under linear constraint under the “spectrum-only” case discussed in Section 5.1; this allows us to design efficient convex programs for convex relaxation of $\mathfrak{R}^{\mathsf{worst}}$.
Definition C.1 (Minimax-optimal pretraining).
The ensuing result provides an justification for Definition C.1, showing that the objective $\mathfrak{R}^{\mathsf{worst}}$, minimized to determine $\widehat{\mathbf{B}}$, tends to approximate the actual worst-case risk.
Proposition C.2.
Proof of Proposition C.2.
Note that
from Theorem 3.3 implies the following
Appendix D Convexity of Spectrum-Only Optimization and Convex Programs
$\hat{\mathbf{q}}_{i}=\mathbf{u}_{i},\forall i\in\{1,\ldots,p\}.$ (72)
which are defined in Section 4 and Appendix C respectively. In this section, we cast the above optimization problems in a more general form
where $f$ is arbitrary function. We show in Section D.1 below that to solve the above, we may solve a linearly constraint optimization problem
where $\tilde{\mathcal{V}},\tilde{\mathfrak{B}}^{\mathsf{avg}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}$ are objectives after applying certain change of variables to $\mathcal{V},\mathfrak{B}^{\mathsf{avg}}$ and $\mathfrak{B}^{\mathsf{avg}}$. We give their definition in Section D.1. A convenient fact that we prove in Section D.2 is that $\tilde{\mathcal{V}},\tilde{\mathfrak{B}}^{\mathsf{avg}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}$ are in in fact convex functions of $\mathbf{x}_{H}$. The above results are important for proving Theorem 5.1. As a byproduct, we also provide convex relaxation of the objectives $\mathfrak{R}^{\mathsf{avg}}$ and
D.1 Equivalence
Given $x_{i}\in[0,1]$ for $i\in\{1,\ldots,p\}$, we define a new objective
where $\mathbf{x}_{H}\in[0,1]^{h}$ with entries $x_{i},i\in H$ and
Proposition D.1 (Equivalence).
Given (14), the solution $\{r_{i}\}_{i=1}^{p}$ of the optimization problem
Proof of Proposition D.1.
Given (14), we have that
where $b_{0}$ is defined as the unique solution of the following equation
$\tilde{\mathcal{F}}\leq\mathcal{F}.$ (80)
D.2 Convexity
Proposition D.2 (Convexity).
Proof of Proposition D.2.
We first note that
$\quantity{\mathbf{x}_{H}\in\mathbb{R}^{h}:h^{-1}\sum_{i\in H}x_{i}=1-n/h}.$
D.3 Algorithmic Implications
The results from Section D.1 and Section D.2 may be leveraged to construct efficient convex algorithms. Using Section D.1 results, we see that to optimize
we only need to solve the following optimization problems
We showed in Section D.2 that $\tilde{\mathcal{V}},\tilde{\mathfrak{B}}^{\mathsf{avg}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}$ are convex by themselves. However, due the presence of the interaction terms $\tilde{\mathfrak{B}}^{\mathsf{avg}}\cdot\tilde{\mathcal{V}}$ and $\tilde{\mathfrak{B}}^{\mathsf{worst}}\cdot\tilde{\mathcal{V}}$, it is uncertain that the objective will remain convex. That said, the objective is in a simple enough form where most constrained local opitmization solvers may be applied. For example, we found that it is quite efficient to solve (81) with the sequential quadratic programming (SQP) routine implemented in Python SciPy package [63].
Therefore, we may consider the following optimization problems
It is easy to verify using Proposition D.2 that the optimization problems (82) are convex program, which can be solved efficiently (see [10] for a review). The solution, obtained from $\hat{r}_{i}\leftarrow c\eta_{i}\quantity(\frac{x_{i}}{1-x_{i}}),i\in H$ will minimize the relaxed upper bounds, $\mathfrak{R}^{\mathsf{avg}}_{\mathsf{rel}}$ and $\mathfrak{R}^{\mathsf{worst}}_{\mathsf{rel}}$, of the original objectives, $\mathfrak{R}^{\mathsf{avg}}$ and $\mathfrak{R}^{\mathsf{worst}}$.
Appendix E Obtain Fully-Optimized $\widehat{\mathbf{B}}$ via Backpropogation
@@
Backpropagating through Eigendecomposition.
@@
Appendix F Upstream Sample Complexity.
In this section, we relax the assumption of oracle knowledge of the ground-truth representation $\mathbf{B}^{\star}$. There are many possible models for how these unknowns could be estimated from pretraining data, and our previous results are not tied to any particular setup. We consider the simple example given in Example F.1 and track how these estimation errors affect the downstream risk.
Example F.1.
In the upstream, the model receives training data for $q$ distinct upstream tasks $\quantity(\mathbf{y}_{\mathsf{pre}}^{(i)},\mathbf{X}_{\mathsf{pre}}^{(i)})_{i=1}^{q}$ where
@@
We recall the setting here for reader’s convenience. Recall that in the upstream, the model receives training data for $q$ distinct upstream tasks $\quantity(\mathbf{y}_{\mathsf{pre}}^{(i)},\mathbf{X}_{\mathsf{pre}}^{(i)})_{i=1}^{q}$ where
Assume that the pretraining data are abundant $n_{\mathsf{pre}}>p$ and the OLS estimators $\tilde{\mathbf{B}}^{\star}=[\tilde{\mathbf{b}}^{(1)},...,\tilde{\mathbf{b}}^{(q)}],\tilde{\mathbf{b}}^{(i)}=\quantity({\mathbf{X}_{\mathsf{pre}}^{(i)}}^{\top}\mathbf{X}_{\mathsf{pre}}^{(i)})^{-1}{\mathbf{X}_{\mathsf{pre}}^{(i)}}^{\top}\mathbf{y}_{\mathsf{pre}}^{(i)}$ are used to estimate $\mathbf{B}^{\star}$. Let ~$\tilde{\mathfrak{R}}^{\mathsf{avg}}$ be the objective $\mathfrak{R}^{\mathsf{avg}}$ based on the inaccurate estimate $\tilde{\mathbf{B}}^{\star}$. The result below characterizes the error in approximating $R$ with ~$\tilde{\mathfrak{R}}^{\mathsf{avg}}$. See proof in Appendix F.
Theorem F.2.
@@
We prove Theorem F.2 below. Recall definition of $\tilde{\mathbf{B}}^{\star}$ from LABEL:SectionPrertain. Let us define the following notation
@@
We see from Lemma F.3 that $\mathfrak{B}-\tilde{\mathfrak{B}}$ depends on the quantity $\norm{(\tilde{\mathbf{B}}^{\star}-\mathbf{B}^{\star})\bm{\alpha}^{\star}}_{2}$. The following lemma bounds this quantity using Hanson-Wright inequality, which we state in Lemma B.1.
Lemma F.4.
@@
Now we are ready to prove the main result Theorem F.2.
Proof of Theorem F.2.
Combining Lemma F.3 and Lemma F.4 yields the following statement: conditioned on $\bm{\alpha}^{\star}$, for any $D>0$, there exists a constant $C(D)$ such that with probability at least $1-p^{-D}$,
@@
along with Lemma 6.1 and Lemma 6.2, we have the further upper bound that
@@
Now using the above, along with Lemma 6.1 and Lemma 6.2, we have the following statement: for any $D>0$, there exists a constant $C(D,M)$ such that with probability at least $1-p^{-D}$,
@@
Appendix G Supplementary Figures
G.1 Compare Fine-Grained Risk Components in Figure 2
@@
G.2 Figure 2 with different choices of $q$
@@
G.3 Ablating $q$, SNR and Common Structure in $\mathbf{B}^{\star}$